OpenStudy (zzr0ck3r):

Let \(f:\mathbb{R} \rightarrow \mathbb{R}\) be a non negative Lebesgue measurable function. For any set \(A\in \mathcal{M}\) (\( \mathcal{M}\) is the \(\sigma\)-algebra of Lebesgue measurable sets), define \(\mu(A):=\int_A f d\lambda\). Show that \(\mu\) is also a measure on \(\mathcal{M}\)

OpenStudy (zzr0ck3r):

@eliassaab this will be the last one ever:) until maybe graduate school:)

OpenStudy (anonymous):

It is clear that \[ \mu(A) \ge 0\\ \mu(\emptyset)=0\\ \] We need to show that \( \mu \) is \(\sigma \)-additive.

OpenStudy (anonymous):

Let \( A_n\) be a sequence of pairwise disjoint measurable sets and let \(A=\cup_{n=1}^\infty A_n \). We need to show that \[ \mu (A) =\sum_{n=1}^\infty \mu (A_n) \]

OpenStudy (anonymous):

Let \( B_k=\cup_{n=1}^k A_n\). It is clear that \(B_k\) is increasing, the union of all \( B_k\) is equal to A and \[ \mu(B_k)=\int_{B_k} f d\lambda=\sum_{n=1}^k \int_{A_n} f d\lambda= \sum_{n=1}^k \mu(A_n) \] By a problem that we did before we have that \[ \lim_{k\to\infty}\int_{B_k} f d\lambda=\int_A fd\lambda=\mu(A)=\\ \lim_{k\to \infty } \sum_{n=1}^k \mu(A_n)=\mu(A)\\ \sum_{n=1}^\infty \mu(A_n)=\mu(A) \] Do not worry, you can always ask me problems.